OpenStudy (anonymous):

a^2 2a --- Divided by --- b^3 b^2

OpenStudy (anonymous):

The expression a^-2 b^3 divided by 1^4 b^-2 is equivalent to ------------------------ (a^(-2)*b^(3))/[1^(4)*b^(-2)] ---= (b^3/a^(2))/(1*b^2) --- = (b^(5))/(a^2) ================